Given the numbers below, fill in the cumulative frequency columns. X     f     cf 9...

  1. Given the numbers below, fill in the cumulative frequency columns.

    X     f     cf

    9     1     
    8     2     
    7     4     
    6     5     
    5     7     
    4     5     
    3     5     
    2     4     
    1     2     
Answer #1
x f cf
9 1 1
8 2 1+2=3
7 4 1+2+4=7
6 5 1+2+4+5=12
5 7 1+2+4+5+7=19
4 5 1+2+4+5+7+5=24
3 5 1+2+4+5+7+5+5=29
2 4 1+2+4+5+7+5+5+4=33
1 2 1+2+4+5+7+5+5+4+2=35